OpenStudy (anonymous):

determine if the lines 9x - 2y = 16 and 2x - 9y = -18 are parallel, perpendicular or neither

OpenStudy (anonymous):

i think its neither but i need confirmation

OpenStudy (shubhamsrg):

can you tell me slopes of both lines ?

OpenStudy (anonymous):

correct me if im wrong here but..... 2.5x and 2/9 x

OpenStudy (anonymous):

make the expression into the form of y= mx +c. where c is a constant and m is the slope of the line. So for example 9x-2y=16. 9x-2y=16 2y = 9x-16 y= 9/2 x - 8 9/2 is the slope that line. Go ahead, Find the other.

OpenStudy (shubhamsrg):

@jonnyrules96 you sure about 2.5 ?

OpenStudy (shubhamsrg):

see @tiaph 's work..

OpenStudy (anonymous):

right right right XD gotcha

OpenStudy (shubhamsrg):

since neither their product is -1 nor they are equal in that case,,you're right,,ans is neither..

OpenStudy (anonymous):

awesome thnx!